# Mokav: Execution-driven Differential Testing with LLMs

### v1

@@ 10 Transformed version of P as a function-based program:
4.3. Baselines
We compare Mokav against two baselines: Pynguin (Lukasczyk and Fraser, 2022) and Differential Prompting (DP) (Li et al., 2023).
Pynguin is a traditional test generation tool that generates random unit tests for Python programs. Pynguin employs evolutionary algorithms (Panichella et al., 2017) to generate regression tests that maximize a fitness function, such as branch coverage. The generated tests call methods of a target module with correctly typed arguments and contain assertions about the methods’ return values. The assertions check the return values with simple types, i.e., int, str, bytes, bool, and None as well as collections of simple types. We consider Pynguin in our evaluation as it is a well-documented tool and recognized as state-of-the-art in test generation for Python (Li et al., 2023; Liu et al., 2024b).
The second baseline that we consider is DP (Li et al., 2023). DP is an LLM-based tool that gets a buggy program and generates fault-inducing tests for it. DP works in three steps. First, it asks the LLM to generate multiple fixed versions for the buggy program. Next, it requests the LLM to produce a set of test cases for the buggy version. Finally, it runs the generated tests on the buggy and fixed versions. If the generated tests produce different outputs on buggy and fixed programs, they are reported as fault-inducing tests.
Note that DP is different from Mokav, as it takes only a buggy program and produces fault-inducing tests for it. In contrast, Mokav takes two programs and generates difference exposing tests. Therefore, we cannot directly compare Mokav with DP. Instead, we isolate DP’s last step and call it DPP (standing for DP Plus). Similar to DP, DPP works in one iteration and generates all the requested tests in that one iteration. Moreover, it only provides one version of the program in the prompt and excludes the descriptions as well as the example test and its corresponding output and execution difference. The prompt is taken unmodified from the replication package of DP (Li et al., 2023).
4.4. Protocol RQ1 (effectiveness)
To answer RQ1, we run Mokav, Pynguin, and DPP on QuixBugs and C4DET. We use all the program pairs in C4DET for our experiments (see Section 4.2).
@@ 4.4. Protocol RQ1 (effectiveness)
For each tool T, we run it on Quixbugs and C4DET and measure three metrics. First, we count the number of $\langle P,Q\rangle$ pairs for which T generates a DET, and call it #$\#Success\_Pairs$. This number is the main metric to compare the DET generation effectiveness of the considered tools. To better understand the effect of Mokav’s iterative approach, we measure this main metric for each of Mokav iterations. If Mokav generates DETs for more pairs after its first iteration, it indicates the effectiveness of the iterative approach. Second, we measure #$\#TT$, which is the total number of tests generated by T. #$\#TT$ indicates the scale of our study. Third, we measure $DET\_Pr$, the number of problems in C4DET that for at least one of their corresponding submission pairs $\langle P,Q\rangle$, T generates a DET. A tool that generates DETs for more problems is effective on a more diverse set of programs, indicating strong external validity.
@@
Table 1. Effectiveness of Mokav compared with related work. QuixBugs dataset contains 32 pairs of programs, and C4DET contains 1,535 pairs of programs for 221 problems. #Success_pair is the number of pairs for which a DET is generated, #TT is the number of all generated tests, and #DET_Pr is the number of problems a DET is generated for at least one of the corresponding pairs.
@@
4.5. Protocol for RQ2 (ablation study)
  • To measure the effect of LLM’s accuracy and check if Mokav is well designed to work with various LLMs, we also run Mokav using two other LLMs. First, we use
    CodeLlama-instruct-7b as an open-source LLM. To keep the resource demand of this experiment manageable, we run this experiment with minimum temperature and generate one sample at each iteration. Second, we also run Mokav with gpt-4o-2024-05-13, as one of the most powerful existing LLMs. We use GPT-4o with two configurations: (1) the same configuration as for CodeLlama, to compare all models with each other, and (2) the default configuration used by Mokav, to assess Mokav’s effectiveness when using the most advanced existing LLM.
4.6. Protocol for RQ3 (impact of P and Q characteristics)
To answer RQ3, we assess the relation between the effectiveness of Mokav and three different characteristics of the input to the tool: the number of tokens in P and Q, the Levenshtein distance between P and Q, and the number of tokens in the longest test for program pair in C4DET. Note that C4DET contains multiple tests for each problem in C4DET. We consider the longest test in this experiment as it suggests how complex the input to the program pair can be.
[…]
To evaluate the relation between DET generation effectiveness and the two other input characteristics (distance between P and Q, and number of tokens in the example test), we follow a similar strategy. The only difference is the metric according to which we sort the $\langle P,Q\rangle$ pairs and split the data into ten subsets. A significant correlation between $SSuccess$ and each of the three considered metrics provides valuable information regarding where we should expect Mokav to work in practice. It can also inform us what are the fundamental reasons why LLMs can generate DETs in some cases and fail to do so in other cases.
5. Experimental Results
5.1. Results for RQ1 (effectiveness)
Figure 3. Effectiveness over iteration. Mokav drives the LLM towards finding difference-exposing tests.
@@ 22 Q: 2
Table 2. Ablation study about the effectiveness of different Mokav components.
Config ID LLM Iterations Samples Temperature Description Ex_Test Exec_Data #Success_Pairs
@@
C7 GPT-4o 10 1 0 73.9% (1,135/1,535)
C8 GPT-4o 10 10 1 85.5% (1,313/1,535)
According to Table 1, Mokav generates a DET for 58.9% (905/1,535) of C4DET pairs at its first iteration and for 81.7% (1,535/1,535) of pairs at its last iteration, which is the highest performance ever reported. First, this shows that the iterative approach significantly improves the effectiveness of Mokav. This is also illustrated in Figure 3. Second, this result shows that Mokav outperforms both strong baselines Pynguin and DPP, which generate DETs for 4.9% and 37.3% of program pairs, respectively. Our manual investigation reveals that Mokav outperforms Pynguin and DPP mainly because it better understands the type of arguments the programs take. This confirms the importance of using an example test in the prompt to hint the LLM regarding the type and structure of inputs.
The next column (“#TT”) shows the total number of tests generated by each tool. DPP generates 153,500 tests for 1,535 pairs in C4DET, because it asks for a fixed number of 100 tests for each pair. Mokav generates batches of 10 tests at each iteration only if the previous iterations fail to yield a DET. Consequently, Mokav generates only 15,350 tests in its first iteration and this number gradually increases to 48,930 at its 10th iteration. This shows that Mokav’s iterative approach also contributes to its cost efficiency by generating new tests step-by-step and only when needed.
All the major takeaways from the results on C4DET also hold for QuixBugs. Most importantly, Mokav outperforms both Pynguin and DPP even at its first iteration. Mokav generates a DET for 100% (32/32) of the program pairs in QuixBugs. This demonstrates the perfect effectiveness of Mokav on this widely-used dataset.
{mdframed}
We evaluate DET generation with Mokav, Pynguin, and DPP on 32 program pairs from QuixBugs and 1,535 program pairs from C4DET. The results clearly show that Mokav outperforms both baselines on both datasets. On QuixBugs, Mokav achieves a perfect score by generating DETs for 100% (31/32) of cases. On C4DET, Mokav produces DETs for 81.7% (1,255/1,535) of the program pairs. Our results also demonstrate that Mokav’s effectiveness improves over iterations, starting at 58.9% and improving up to 81.7%. To our knowledge, Mokav is the best performing, state-of-the-art approach and tool for generating DETs.
@@ 11 The output produced by P and Q on the generated DET:
5.2. Results for RQ2 (ablation study)
Table 2 presents the results of our ablation study. It compares the effectiveness of different Mokav configurations in terms of their #Success_Pairs. Eight configurations, C1-C8, are considered, with C1 being the default setup of Mokav. C2, C3, and C4 represent configurations where the description, the example test, and the execution data are excluded from the prompt, respectively. Note that in C3, as the example test is excluded, the execution data that depends on the presence of an example test is excluded as well. In C5, temperature=0 is used and only the top one sample is requested from the model at each iteration. C6 and C7 are similar to C5, but use different LLMs, GPT-4o and CodeLlama, respectively. Finally, C8 has the same configuration as C1, but uses GPT-4o as the LLM. The last column of Table 2 shows the #Success_Pairs, the number of pairs in C4DET for which each configuration generates a DET.
C6, C7, and C8 use different LLMs. The table shows that the model in use influences effectiveness. CodeLlama underperforms with an effectiveness of only 11.4% (C6 $\ll$ C5). Expectedly, the very recent GPT-4o (released in May 2024) surpasses GPT-3.5 (C7 $\gg$ C5).
Finally, C8 shows results for the last experiment, where we run Mokav with default settings similar to C1 using GPT-4o, released while we were preparing this manuscript. The results show that the effectiveness is improved to 85.5%, the highest among all configurations. This improvement compared to C1 shows that using a stronger LLM that better understands the semantics of programs directly impacts effectiveness. We note that C8 has an 11.6% improvement over C7 (85.5% vs. 73.9%). This significant improvement reaffirms the soundness of the default setting used in Mokav: a high sample size (10) and high temperature (10) lead to generating a diverse set of useful test cases.
We now discuss an interesting case in LABEL:lst:gpt4-ex, which shows the code diff between a pair of programs, and input and output. The pair comes from author “60,724” for problem 226B on Codeforces. Only GPT-4.o succeeds in generating a valid DET, with exactly 50 characters as input (“BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBB
BBBBBBBBBBBBBBBBBBG”). This shows that the increasingly more powerful LLMs can be very valuable for generating DETs.
@@ 5.2. Results for RQ2 (ablation study)
Our comparison of 8 different configurations of Mokav shows that using Mokav’s default setup and GPT-4o surpasses all other configurations with an effectiveness of 85.5% (1,313/1,535). Among components in Mokav’s approach, the example test given tothe LLM plays the most important role. Using a high temperature and sample size effectively leverages the LLM’s ability to explore various parts of the input space.
5.3. Results for RQ3 (impact of P and Q characteristics)
Table 3. The correlation between Mokav’s effectiveness and P & Q characteristics.
SRC_Tok Test_Tok Lev_Diff
Med SSuccess Med SSuccess Med SSuccess
S1 74 0.96 1 0.89 0.44 0.92
S2 115 0.88 2 0.84 0.68 0.90
S3 153 0.84 3 0.79 0.80 0.86
S4 187 0.86 4 0.77 0.87 0.76
S5 220 0.77 6 0.77 0.91 0.76
S6 263 0.79 9 0.83 0.94 0.79
S7 312 0.77 13 0.78 0.96 0.77
S8 384 0.81 21 0.68 0.99 0.80
S9 496 0.75 55 0.84 0.99 0.79
S10 809.5 0.72 82 0.92 0.99 0.78
Cor -0.89 -0.01 -0.45
p-val(<0.05) 0.0004 (✓) 0.97 (✗) 0.18 (✗)
@@ 1 The diff between P and Q:
Table 3 shows the results of the RQ3 experiment. In this table, S1-S10 represent the subsets of program pairs after split according to each considered program pair characteristics. The table has three main vertical sections representing each of those three characteristics: “SRC_Tok”, the number of tokens in a program pair; “Test_Tok”, the number of tokens in the longest test in C4DET for a program pair; and “Lev_Diff”, the Levenshtein distance between a pair of programs. For each characteristic, the table shows the median value for each subset (“Med”) and the ratio of pairs in each subset for which Mokav successfully generates a DET (“SSuccess”). For example, consider S1. The median number of tokens in the 10% shortest pairs is 74. Mokav generates a DET for 96% of these program pairs. The median number of tokens in the test for 10% of program pairs that have the shortest test is 1. Mokav generates a DET for 89% of these program pairs. Finally, the median Levenshtein distance for the 10% of program pairs with the shortest distance is 0.44. For 92% of these program pairs, Mokav successfully generates a DET.
@@
The last two rows of Table 3 show the correlation between the subset index and each of the considered characteristics. There is a strong negative correlation of -0.89 between the number of tokens in source code and the success rate of Mokav, meaning that the longer the input programs, the less likely Mokav can generate a valid DET. This correlation is statistically significant as its corresponding p-value is 0.0004, much lower than the typical 0.05 threshold. Overall, Table 3 shows that the number of tokens in program pairs is the only characteristic with a statistically significant correlation with Mokav’s success rate. This indicates that to apply Mokav on multi-function programs, we would have to employ powerful LLMs with large context windows, and delicately provide them with runtime hints regarding program semantics.
@@
The absence of a significant correlation between the length of the longest test in C4DET for a program pair and Mokav’s success rate demonstrates the following: the size of the inputs for a program pair does not directly tell how challenging it is to find the corner cases that expose differences. We observe no significant correlation between the program pair’s distance and Mokav’s success rate. At first glance, it is easier to expose the difference between two programs if they are textually distant; therefore, we expect to see a positive and significant correlation between Mokav’s success rate and the distance between pairs. On the other hand, we notice that in C4DET, program pairs close to each other tend to be small and easier to understand. This means we expect the LLM can better understand program pairs with small distances and thus generate a valid DET for them. Consequently, the distance between program pairs impacts Mokav’s success rate from two opposite directions. This leads to the absence of a significant correlation between Mokav’s effectiveness and distance between P&Q.
{mdframed}
@@
Our experiments show a significant correlation between the length of program pairs and Mokav’s success rate in generating a DET for them. There is strong statistical evidence that LLMs better capture the semantics of small program pairs to find corner cases that expose their differences. We do not find a significant correlation between Mokav’s effectiveness and the distance between program pairs.
6. Threats to Validity

### v2

@@ 10 Transformed version of P as a function-based program:
4.3 Baselines
We compare Mokav against two baselines: Pynguin [16] and Differential Prompting (DP) [17].
Pynguin is a traditional test generation tool that generates random unit tests for Python programs. Pynguin employs evolutionary algorithms [27] to generate regression tests that maximize a fitness function, such as branch coverage. The generated tests call methods of a target module with correctly typed arguments and contain assertions about the methods’ return values. The assertions check the return values with simple types, i.e., int, str, bytes, bool, and None as well as collections of simple types. We consider Pynguin in our evaluation as it is a well-documented tool and recognized as state-of-the-art in test generation for Python [17, 25].
The second baseline that we consider is DP [17]. DP is an LLM-based tool that gets a buggy program and generates fault-inducing tests for it. DP works in three steps. First, it asks the LLM to generate multiple fixed versions for the buggy program. Next, it requests the LLM to produce a set of test cases for the buggy version. Finally, it runs the generated tests on the buggy and fixed versions. If the generated tests produce different outputs on buggy and fixed programs, they are reported as fault-inducing tests.
Note that DP differs from Mokav in two major perspectives. First, the main goal of DP is to generate a test that detects the fault in a single given buggy program. In contrast, Mokav aims to generate a test that exposes functional differences between two different programs. This key difference between the goals of Mokav and DP leads to various differences between their prompting strategies. Namely, DP prompts only contain information about the buggy program, while Mokav includes information about both versions of the program. Second, Mokav adopts an iterative approach that directs the LLM toward DET generation with execution-based feedback, while DP calls the LLM only once and does not provide execution data in the prompt. Given these major differences, we cannot directly compare Mokav with DP. Instead, we isolate DP’s last two steps and call it DPP (standing for DP-Prime). Similar to DP, DPP works in one pass and generates all the requested tests in that one single iteration. The prompt is taken unmodified from the replication package of DP [17].
4.4 Protocol RQ1 (effectiveness)
To answer RQ1, we run Mokav, Pynguin, and DPP on QuixBugs and C4DET. We use the transformed version of all program pairs in C4DET for our experiments (see Section 4.2).
@@ 4.4 Protocol RQ1 (effectiveness)
[…]
@@
4.5 Protocol for RQ2 (ablation study)
  • To measure the effect of LLM’s accuracy and check if Mokav is well designed to work with various LLMs, we also run Mokav using three other LLMs. First, we use two open-source LLMs from two different enterprises: CodeLlama-instruct-7b by Meta and Codestral-2501 by Mistral AI. Codestral is one of the most recent LLMs as of April 2025; it is a lightweight, low-latency model designed for code and reasoning tasks. This makes Codestral an ideal LLM for evaluating how far we can push the effectiveness of Mokav using affordable models. To keep the resource demand of this experiment manageable, we run this experiment with the minimum temperature and generate one sample at each iteration. In addition to the open-source models, we also run Mokav with gpt-4o-2024-05-13, as one of the most powerful existing LLMs. We use GPT-4o with two configurations: (1) the same configuration as for CodeLlama and Codestral, to compare all models with each other, and (2) the default configuration used by Mokav, to assess Mokav’s effectiveness when using the most advanced existing LLM.
4.6 Protocol for RQ3 (impact of P and Q characteristics)
To answer RQ3, we assess the relation between the effectiveness of Mokav and four different characteristics of the input to the tool: the total number of tokens in P and Q, the cyclomatic complexity of Q, the Levenshtein distance between P and Q, and the number of tokens in the longest test for program pair in C4DET. Note that C4DET contains multiple tests for each problem in C4DET. We consider the longest test in this experiment as it suggests how complex the input to the program pair can be.
[…]
To evaluate the relation between DET generation effectiveness and the three other input characteristics (cyclomatic complexity of Q, distance between P and Q, and number of tokens in the example test), we follow a similar strategy. The only difference is the metric according to which we sort the $\langle P,Q\rangle$ pairs and split the data into ten subsets. A significant correlation between $SSuccess$ and each of the four considered metrics provides valuable information regarding where we should expect Mokav to work in practice. It can also inform us of the fundamental reasons why LLMs can generate DETs in some cases and fail to do so in other cases.
5 Experimental Results
5.1 Results for RQ1 (effectiveness)
Table 1: Effectiveness of Mokav compared with related work. QuixBugs dataset contains 32 pairs of programs, and C4DET contains 1,535 pairs of programs for 221 problems. #Success_pair is the number of pairs for which a DET is generated, #TT is the number of all generated tests, and #DET_Pr is the number of problems a DET is generated for at least one of the corresponding pairs.
@@ 5.1 Results for RQ1 (effectiveness)
Figure 3: Effectiveness over iteration. Mokav drives the LLM towards finding difference-exposing tests.
@@ 22 Q: 2
According to Table 1, Mokav generates a DET for 58.9% (905/1,535) of C4DET pairs at its first iteration and for 81.7% (1,535/1,535) of pairs at its last iteration, which is the highest performance ever reported. First, this shows that the iterative approach significantly improves the effectiveness of Mokav. This is also illustrated in Figure 3. Second, this result shows that even at its initial attempt, Mokav outperforms both strong baselines Pynguin and DPP:  toolname’s initial attempt generates a DET for 58.9% of pairs, while Pynguin and DPP generate DETs for 4.9% and 37.3% of program pairs, respectively. While the iteration phase significantly improves Mokav effectiveness, Mokav achieves strong DET generation results even without paying the cost of iterative LLM invocations. Our manual investigation reveals that Mokav outperforms Pynguin and DPP mainly because it better understands the type of arguments the programs take. This confirms the importance of using an example test in the prompt to hint to the LLM regarding the type and structure of inputs.
The next column (“#TT”) shows the total number of tests generated by each tool. DPP generates 153,500 tests for 1,535 pairs in C4DET, because it asks for a fixed number of 100 tests for each pair. Mokav generates batches of 10 tests at each iteration only if the previous iterations fail to yield a DET. Consequently, Mokav generates only 15,350 tests in its first iteration, and this number gradually increases to 48,930 at its 10th iteration. This shows that Mokav’s iterative approach also contributes to its cost efficiency by generating new tests step-by-step and only when needed.
All the major takeaways from the results on C4DET also hold for QuixBugs. Most importantly, Mokav outperforms both Pynguin and DPP even at its first iteration. Mokav generates a DET for 100% (32/32) of the program pairs in QuixBugs. This demonstrates the perfect effectiveness of Mokav on this widely-used dataset. We further analyze our experimental results to understand the reasons behind this perfect effectiveness. We first notice that the effectiveness of Mokav cannot be attributed to data leakage in the LLM training dataset, as DPP also uses the same LLM for test generation and fails in 50% (16/32) of the cases. We discuss the threats of data leakage in Section 6. Our analysis reveals that QuixBugs DETs are close to the example tests in the dataset. Therefore, adding the example test to Mokav prompts significantly helps the LLM in generating DETs. This finding explains the difference between the effectiveness of DPP and Mokav, which lies in the provided example test.
We evaluate Mokav, Pynguin, and DPP for generating difference exposing tests on 32 program pairs from QuixBugs and 1,535 program pairs from C4DET. The results clearly show that Mokav outperforms the state of the art on both datasets. On QuixBugs, Mokav achieves a perfect score by generating DETs for 100% (31/32) of cases. On C4DET, Mokav produces DETs for 81.7% (1,255/1,535) of the program pairs. Our results also demonstrate that Mokav’s effectiveness improves over iterations, starting at 58.9% and improving up to 81.7%. To our knowledge, Mokav is the best performing approach and tool for generating DETs.
5.2 Results for RQ2 (ablation study)
Table 2 presents the results of our ablation study. It compares the effectiveness of different Mokav configurations in terms of their #Success_Pairs. Nine configurations, C1-C9, are considered, with C1 being the default setup of Mokav. C2, C3, and C4 represent configurations where the description, the example test, and the execution data are excluded from the prompt, respectively. Note that in C3, as the example test is excluded, the execution data that depends on the presence of an example test is excluded as well. In C5, temperature=0 is used and only the top one sample is requested from the model at each iteration. C6, C7 and C8 are similar to C5, but use different LLMs, CodeLlama, Codestral, and GPT-4o, respectively. Finally, C9 has the same configuration as C1, but uses GPT-4o as the LLM. The last column of Table 2 shows the #Success_Pairs, the number of pairs in C4DET for which each configuration generates a DET.
Table 2: Ablation study about the effectiveness of different Mokav components.
Config ID LLM Iter. Samp. Temp. Desc. Ex_Test Exec_Data #Success_Pairs
@@ 5.2 Results for RQ2 (ablation study)
C7 Codestral 10 1 0 64.4% (989/1,535)
C8 GPT-4o 10 1 0 73.9% (1,135/1,535)
C9 GPT-4o 10 10 1 85.5% (1,313/1,535)
@@ 11 The output produced by P and Q on the generated DET:
C5, C6, C7, and C8 use different LLMs. The table shows that the model in use impacts effectiveness. CodeLlama underperforms our default GPT-3.5 with an effectiveness of only 11.4% (C6 $\ll$ C5). In contrast, Codestral outperforms GPT-3.5 (C7 $\gg$ C5) by a large margin, succeeding on 64.4% (989/1,535) of the program pairs. We note that the total cost of DET generation with Codestral for 1,535 pairs in our dataset was 6.45 USD. This indicates the practicality of Mokav, as it achieves strong effectiveness by leveraging affordable open-source models that are optimized for coding tasks. Also, as expected, the highly strong and quite expensive GPT-4o surpasses GPT-3.5 (C8 $\gg$ C5), CodeLlama (C8 $\gg$ C6), and Codestral (C8 $\gg$ C7). This demonstrates Mokav’s potential to harness the power of frontier models for DET generation.
Given the low cost of open-source models, such as Codestral, we can increase the sample size at each iteration while keeping the test generation cost reasonable. As our comparison between C1 and C5 shows, generating more samples can improve the effectiveness of Mokav. Therefore, we expect to achieve greater effectiveness by increasing the Codestral sample size, even beyond the Mokav default configuration of $sample\_size=10$. This technique can be evaluated in the future to see if we can replace expensive closed-source models with affordable transparent open-source ones without losing effectiveness.
Finally, C9 shows the results for the last experiment, where we run Mokav with default settings similar to C1 using GPT-4o, the best model in our experiments. The results show that effectiveness is improved to 85.5%, the highest among all configurations. This improvement compared to C1 shows that using a stronger LLM that better understands the semantics of programs directly impacts effectiveness. This significant improvement reaffirms the soundness of the default setting used in Mokav: a high sample size (10) and high temperature (10) lead to generating a diverse set of useful test cases.
We now discuss an interesting case in LABEL:lst:gpt4-ex, which shows the code diff between a pair of programs, and input and output. The pair comes from author “60,724” for problem 226B on Codeforces. Among all the configs in Table 2, only C8, which uses GPT-4.o with temperature=0 and ten iterations of execution-based feedback, succeeds in generating a valid DET. Our analysis also shows that DPP and Pynguin fail to generate a valid DET in this case. The valid DET generated by C8 contains exactly 50 characters as input (“BBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBBB
BBBBBBBG”). This example demonstrates two facts. First, the increasingly more powerful LLMs are valuable for generating DETs, as they better understand the code semantics. Secondly, this example shows that in out-of-distribution cases the execution-based feedback is needed on best LLM solutions for guiding the LLM toward DET generation.
@@
Our comparison of nine different configurations of Mokav shows that using Mokav’s optimized setup with a frontier model (GPT-4o) surpasses all other configurations with a maximum effectiveness of 85.5% (1,313/1,535). Among components in Mokav’s approach, the example test given to the LLM plays the most important role. Using a high temperature and sample size effectively leverages the LLM’s ability to explore various parts of the input space.
5.3 Results for RQ3 (impact of P and Q characteristics)
Listing 5: A pair for which no config generates a DET.
@@ 1 The diff between P and Q:
Table 3 shows the results of the RQ3 experiment. In this table, S1-S10 represent the subsets of program pairs after split according to each considered program pair characteristics. The table has four main vertical sections representing each of those four characteristics: “SRC_Tok”, the number of tokens in a program pair; “Cyclo_Comp”, the cyclomatic complexity of Q, which represents the correct version of the program pair in C4DET; “Test_Tok”, the number of tokens in the longest test in C4DET for a program pair; and “Lev_Diff”, the Levenshtein distance between a pair of programs. For each characteristic, the table shows the median value for each subset (“Med”) and the ratio of pairs in each subset for which Mokav successfully generates a DET (“SSuccess”). For example, consider S1. The median number of tokens in the 10% shortest pairs is 74. Mokav generates a DET for 96% of these program pairs. The median cyclomatic complexity of the 10% least complex programs is 1, and Mokav generates a DET for 91% of these program pairs. The median number of tokens in the test for 10% of program pairs that have the shortest test is 1. Mokav generates a DET for 89% of these program pairs. Finally, the median Levenshtein distance for the 10% of program pairs with the shortest distance is 0.44. For 92% of these program pairs, Mokav successfully generates a DET.
Table 3: The correlation between Mokav’s effectiveness and P & Q characteristics.
SRC_Tok Cyclo_Comp Test_Tok Lev_Diff
Med SSuccess Med SSuccess Med SSuccess Med SSuccess
S1 74 0.96 1 0.91 1 0.89 0.44 0.92
S2 115 0.88 2 0.86 2 0.84 0.68 0.90
S3 153 0.84 3 0.80 3 0.79 0.80 0.86
S4 187 0.86 4 0.80 4 0.77 0.87 0.76
S5 220 0.77 4 0.80 6 0.77 0.91 0.76
S6 263 0.79 5 0.79 9 0.83 0.94 0.79
S7 312 0.77 6 0.83 13 0.78 0.96 0.77
S8 384 0.81 7 0.78 21 0.68 0.99 0.80
S9 496 0.75 9 0.81 55 0.84 0.99 0.79
S10 809.5 0.72 14 0.72 82 0.92 0.99 0.78
Cor -0.89 -0.63 -0.01 -0.45
p-val(<0.05) 0.0004 (✓) 0.04 (✓) 0.97 (✗) 0.18 (✗)
@@
The last two rows of Table 3 show the correlation between the subset index and each of the considered characteristics. There is a strong negative correlation of -0.89 between the number of tokens in the source code and the success rate of Mokav. Similarly, there is a strong negative correlation of -0.63 between the cyclomatic complexity of the program and the success rate of Mokav. Both of these correlations are statistically significant with a p-value of 0.0004 and 0.04, respectively. These statistically significant negative correlations highlight that Mokav is less likely to generate a valid DET for more syntactically complex program pairs. In sum, Table 3 shows that the length and complexity of program pairs are the only characteristics with a statistically significant correlation with the success rate of Mokav. This indicates that to apply Mokav on lengthy and complex multi-function programs, we would have to employ powerful LLMs with large context windows, and delicately provide them with runtime hints regarding program semantics.
@@
The absence of a significant correlation between the length of the longest test in C4DET for a program pair and Mokav’s success rate demonstrates the following: the size of the inputs for a program pair does not directly tell how challenging it is to find the corner cases that expose differences. We observe no significant correlation between the program pair’s distance and Mokav’s success rate. At first glance, it is easier to expose the difference between two programs if they are textually distant; therefore, we expect to see a positive and significant correlation between Mokav’s success rate and the distance between pairs. On the other hand, we notice that in C4DET, program pairs close to each other tend to be smaller and easier to understand. This means we expect the LLM to better understand program pairs with small distances and thus generate a valid DET for them. Consequently, the distance between program pairs impacts Mokav’s success rate from two opposite directions. This leads to the absence of a significant correlation between Mokav’s effectiveness and distance between P&Q.
@@
Our experiments show a significant correlation between the length and cyclomatic complexity of program pairs and Mokav’s success rate in generating a difference exposing test. There is strong statistical evidence that LLMs better capture the semantics of syntactically simple program pairs to find corner cases that expose their differences. We do not find a significant correlation between Mokav’s effectiveness and the distance between program pairs.
6 Discussion
6.1 Threats to Construct Validity